# Phantom Edges in the Problem Hamiltonian: A Method for Increasing Performance and Graph Visibility for QAOA

### v1

@@ Abstract
The Quantum Approximate Optimization Algorithm (QAOA) is a variational quantum algorithm designed to solve combinatorial optimization problems. However, a key limitation of QAOA is that it is a “local algorithm,” meaning it can only optimize over local properties of the graph for finite circuit depths. In this work, we present a new QAOA ansatz that introduces only one additional parameter to the standard ansatz—regardless of system size—allowing QAOA to “see” more of the graph at a given depth $p$. We achieve this by modifying the target graph to include additional $\alpha$-weighted edges, with $\alpha$ serving as a tunable parameter. This modified graph is then used to construct the phase operator and allows QAOA to explore a wider range of the graph’s features for a smaller $p$. We derive a general formula for our new ansatz at $p=1$ and analytically show an improvement in the approximation ratio for cycle graphs. We also provide numerical experiments that demonstrate significant improvements in the approximation ratio for the Max-Cut problem over the standard QAOA ansatz for $p=1$ and $p=2$ on random regular graphs up to 16 nodes.
@@ I Introduction
Over the past several years, researchers have studied the potential ability of quantum computers to outperform classical computers on a variety of problems and tasks. One such category of problems are combinatorial optimization problems, where – given a finite discrete set of candidate solutions – the objective is to find an optimal solution with respect to some cost function. Such problems appear in many different domains including power systems [19], logistics [16], VLSI [2], and more. Our work focuses on the well-studied Max-Cut problem which is to determine an optimal partition of the vertices of the graph into two groups such that the number of edges between the two groups is maximized; however, the main ideas and techniques in our work can also be used for other combinatorial optimization problems.
@@
Since the cost Hamiltonian typically encodes the cost function of the underlying combinatorial optimization problem, it is not clear a priori what alternatives one should consider. In particular, with the exception of an affine shift in the cost function, any other cost Hamiltonian will most likely break QAOA’s connection to Quantum Adiabatic Computing (QAC); this connection ensures that QAOA (with optimal parameters) approaches an optimal solution as the number of circuit layers $p$ tends to infinity. In this work, we consider an alternative cost Hamiltonian that is formed by a modified graph which is constructed by starting with the original input graph and adding additional $\alpha$-weighted edges, with $\alpha$ being a tunable parameter. If one allows optimization of the $\alpha$-parameter, then this QAOA variant does at least as well as standard QAOA, as the standard QAOA is recovered at $\alpha=0$.
@@
Similar to Multi-Angle QAOA, our method introduces additional parameters; however, our method only adds a single parameter, regardless of problem size or number of QAOA layers. On the other hand, Multi-Angle QAOA adds an additional $p(n+m-2)$ parameters where $m$ is the number of terms in the original cost Hamiltonian; the addition of such parameters significantly increases the parameter search space, thus making the cost landscape increasingly difficult and computationally intensive to search over.
In this work, we demonstrate theoretical results for single-layer QAOA with our approach. For any way of adding $\alpha$-weighted edges, we provide a general formula for the expectation of an edge being cut for single-layer QAOA. Using this formula, we demonstrate an example where our approach does strictly better than standard QAOA, more specifically, for even-lengthed cycles, we show that, when placing the additional edges in $G$ to form triangles, our ansatz yields an approximation ratio of $0.7925$ whereas standard QAOA only has an approximation ratio of $0.75$.
@@ II Background
II-A Quantum Approximate Optimization Algorithm
QAOA is a hybrid quantum-classical algorithm where a parameterized quantum circuit generates sample solutions, and a classical optimizer adjusts the circuit parameters to maximize (or minimize) the objective function. The QAOA ansatz is inspired by adiabatic quantum computing [1], alternating between the mixer operator $U(B,\beta_{i})$ and the phase operator $U(C,\gamma_{i})$, where $\beta_{i}$ and $\gamma_{i}$ are the respective parameters for the $i$th layer. The number of layers, $p$, is referred to as the circuit depth.
@@ II-A Quantum Approximate Optimization Algorithm
The phase operator typically encodes the problem we wish to solve. For the MaxCut problem, the phase operator is generated by the cost Hamiltonian $C$ that represents the cut value of a given partition of the graph’s vertices. The phase operator is therefore:
@@
The variational quantum state after $p$ layers is given by
@@
where |$\ket{s}$ is the initial state. In this paper, we follow the standard practice of initializing the state in the most excited state of the mixer Hamiltonian: |$\ket{+}^{\otimes n}$. Note that this initial state is the equal superposition of all $2^{n}$ possible cuts in the graph. Equivalently, measuring |$\ket{s}$ would cause each vertex to independently have equal probability of being on either side of the cut; one can show that this implies that each edge is cut with probability $1/2$ and hence measuring |$\ket{s}$ would yield a cut value of $1/2$ the number of edges.
@@
Maximizing the expectation value of $\langle C\rangle$ with respect to the parameters $\boldsymbol{\gamma}$, $\boldsymbol{\beta}$ is the objective of QAOA. Let $F_{p}$ represent the expectation value of the cost Hamiltonian after $p$ layers. The maximum achievable expectation value, $\text{max}_{\boldsymbol{\gamma},\boldsymbol{\beta}}F_{p}(\gamma,\beta)$, increases with the circuit depth $p$. In the limit as $p\to\infty$, QAOA (with optimal parameters) is guaranteed to achieve an approximation ratio of $1$ [10].
II-B Graph Visibility in QAOA
@@
To illustrate QAOA’s locality, let’s evaluate a single edge term $C_{uv}=\frac{1}{2}(I-Z_{u}Z_{v})$ in the cost Hamiltonian (recall that $C=\sum_{uv}C_{uv}$). The expected cut value of the edge $(u,v)$ at depth $p=1$ is
@@
The only terms in the mixer Hamiltonian $B=\sum_{j=1}^{n}X_{j}$ that will contribute to the expected cost are $X_{u}$ and $X_{v}$, as the rest will commute through and cancel. Similarly, the only terms that remain from the cost Hamiltonian $C=\sum_{uv}C_{uv}$ will correspond to the edges that involve vertices $u$ or $v$. Therefore, when maximizing the total cost Hamiltonain $C$, we are maximizing over the linear combination of subgraphs for which edges are within single hop from either $u$ or $v$. This analysis can easily be extended to depth $p$, for which it can be shown that the subgraphs will be composed of edges that are $p$-hops away from $u$ or $v$.
However, adjusting the cost Hamiltonian that is used in the phase operatory $U(C,\gamma)$ can lead to changes in how much the QAOA can ”see” for a given $p$. One example of the application of a modified phase operator is shown in [22]. We discuss our modified ansatz in detail in the next section.
III Modified Phase Operator
Mathematically, we introduce a modified phase operator $U(\tilde{C},\gamma)$, where $\tilde{C}$ is the Hamiltonian including additional small-weight edges:
@@ III Modified Phase Operator
Here, $C$ is the original cost Hamiltonian, $C^{\prime}$ represents the additional edge terms, and $\alpha$ is the weight of the additional edges. The modified phase operator is then:
@@
III-A Constructing the Modified Graph
In this paper, we explore two methods for constructing the modified graph $G^{\prime}$, referred to as the ”Full Method” and the ”Triangle Method.”
The Full Method involves adding $\alpha$-weighted edges between all pairs of vertices in $G$ that do not already share an edge, effectively creating a fully connected graph $G^{\prime}$, in other words the set of edges that are added are simply the edges of the complement of $G$. Mathematically, if $E$ is the set of edges in the original graph $G$, the new set of edges in $G^{\prime}$, denoted by $E^{\prime}$, includes all possible edges between vertices:
@@ III-A Constructing the Modified Graph
The additional edges in $G^{\prime}$ are then weighted by the factor $\alpha$ in the modified cost Hamiltonian $C^{\prime}$.
The Triangle Method is a more strategic approach that adds $\alpha$-weighted edges only between vertices that are two hops away in $G$, forming triangles with the existing edges. Mathematically, for any vertex pair $(u,v)$ such that the shortest path between $u$ and $v$ in $G$ is exactly two edges, we add an edge $(u,v)$ to $G^{\prime}$:
@@
IV Derivation of Max-Cut Expectation Value with $G^{\prime}$ Phase Operator
We derive a formula for our modified ansatz at $p=1$, where each edge added to $G$ to construct $G^{\prime}$, as encoded in the modified phase operator, is assigned a weight of $\alpha$. The derivation and resulting equation are similar to Eq. 5.10 in [11]. In this section, we present the theorem and its proof for our ansatz to ensure completeness. We acknowledge that the expression itself may not provide immediate insights or be easily interpretable. Therefore, we provide an example to improve interpretability and demonstrate the effectiveness of introducing additional edges such that triangles are formed with the edges in $G$.
@@ IV Derivation of Max-Cut Expectation Value with $G^{\prime}$ Phase Operator
Let $G$ be the target optimization graph and let $G^{\prime}$ be the graph that includes additional edges weighted by $\alpha$. Let’s define $N$ as the set of neighbors in $G$, and $N^{\prime}$ as the set of neighbors in $G^{\prime}$ but not in $G$. For each edge $(uv)$ in the edge set $E$ of $G$, the expectation value $\langle C_{uv}\rangle$ at depth $p=1$ is given by
$\langle C_{uv}\rangle$ $=\frac{1}{2}+\frac{\sin(4\beta)}{4}\sin\gamma\left(\cos^{d}\gamma\cos^{d^{\prime}}\alpha\gamma+\cos^{e}\gamma\cos^{e^{\prime}}\alpha\gamma\right)$
@@
$\quad\quad(\cos(\gamma)\cos(\alpha\gamma)+\sin(\gamma)\sin(\alpha\gamma))^{f^{\prime}}$
$\quad\quad-(\cos(\gamma)\cos(\alpha\gamma)-\sin(\gamma)\sin(\alpha\gamma))^{f^{\prime}}\bigg{)}$
@@
$\langle\gamma,\beta|C_{uv}|\gamma,\beta\rangle=\frac{1}{2}-\frac{1}{2}\langle s|e^{i\gamma\tilde{C}}e^{i\beta B}Z_{u}Z_{v}e^{-i\beta B}e^{-i\gamma\tilde{C}}|s\rangle.$
@@
$\langle s|e^{i\gamma\tilde{C}}\left(c^{2}Z_{u}Z_{v}+sc(Y_{u}Z_{v}+Z_{u}Y_{v})+s^{2}Y_{u}Y_{v}\right)e^{-i\gamma\tilde{C}}|s\rangle.$ (3)
@@
$=\langle s|(\cos(\gamma)I-i\sin(\gamma)Z_{u}Z_{v})\prod_{i=1}^{d}(\cos(\gamma)I-i\sin(\gamma)Z_{u}Z_{c_{i}})$
×$\quad\times\prod_{i=1}^{d^{\prime}}(\cos(\alpha\gamma)I-i\sin(\alpha\gamma)Z_{u}Z_{c_{i}^{\prime}})|s\rangle,$ (4)
where $d=|N(u)\setminus v|$, $c_{i}$ is a vertex in $N(u)\setminus v$ such that $(u,c_{i})\in E$, $d^{\prime}=|N^{\prime}(u)|$, and $c^{\prime}_{i}$ is a vertex in $N^{\prime}(u)$ such that $(u,c^{\prime}_{i})\in E^{\prime}$. The only term that contributes to the expectation value in Eq. IV is the one proportional to $Z_{u}Z_{v}\otimes I^{\otimes d}\otimes I^{\otimes d^{\prime}}\cdot Y_{u}Z_{v}=-iX_{u}$. Thus, we have:
@@
Due to symmetry, the expression for the expectation value of the term $Z_{u}Y_{v}$ will be the same as in Eq. IV, except now we use the variables $e=|N(v)\setminus u|$ connected by an edge in $E$ and $e^{\prime}=|N^{\prime}(v)|$ connected by an edge in $E^{\prime}$. Thus,
@@
$=\sin\gamma\left(\cos^{d}\gamma\cos^{d^{\prime}}\alpha\gamma+\cos^{e}\gamma\cos^{e^{\prime}}\alpha\gamma\right),$ (6)
For the final term $Y_{u}Y_{z}$, we follow a similar procedure as detailed in [11] except we inlcude the weighted terms $\alpha Z_{u}Z_{wi}$ from $\tilde{C}$:
$\langle s|e^{i\gamma C}Y_{u}Y_{v}e^{-i\gamma C}|s\rangle=\langle s|\prod_{i=1}^{d}(\cos(\gamma)I-i\sin(\gamma)Z_{u}Z_{w_{i}})$
@@
×$\quad\times\prod_{j=1}^{e^{\prime}}(\cos(\alpha\gamma)I-i\sin(\alpha\gamma)Z_{v}Z_{w^{\prime}_{j}})Y_{u}Y_{v}|s\rangle.$
The analysis is similar to the one detailed in [11] for deriving Eq. 5.14, except now we must consider the following three cases for a pair of vertices $(u,v)\in E$:
@@
    Both edges $(u,w_{i})$ and $(v,w_{i})$ belong to the original graph $G$.
@@
    One of the edges, either $(u,w_{i})$ or $(v,w_{i})$, belongs to the original graph $G$, while the other belongs to the modified graph $G^{\prime}$.
@@
    Both edges $(u,w_{i})$ and $(v,w_{i})$ belong to the modified graph $G^{\prime}$.
Therefore,
@@
where $f$, $f^{\prime}$, and $f^{\prime\prime}$ represent the number of triangles in cases 1, 2, and 3, respectively. Using the binomial theorem and trigonometric identities, we simplify the expression to:
@@
$=\cos^{d+e-2f}c^{\prime\prime d^{\prime}+e^{\prime}}(1-\cos^{f}(2\gamma))$
$+\cos^{d+e-f^{\prime}}c^{\prime\prime d^{\prime}+e^{\prime}-f^{\prime}}\frac{1}{2}\left((c^{\prime}c^{\prime\prime}+s^{\prime}s^{\prime\prime})^{f^{\prime}}-(c^{\prime}c^{\prime\prime}-s^{\prime}s^{\prime\prime})^{f^{\prime}}\right)$
$+\cos^{d+e}c^{\prime\prime d^{\prime}+e^{\prime}-2f^{\prime\prime}}(1-\cos^{f^{\prime\prime}}(2\alpha\gamma)).$ (7)
Substituting Eq.  IV and Eq.  IV into Eq.  3 we obtain the final expression for $\langle C_{uv}\rangle$ in terms of $\gamma,\beta$ and $\alpha$.

### v2

@@ Abstract
The Quantum Approximate Optimization Algorithm (QAOA) is a variational quantum algorithm that can be used to approximately solve combinatorial optimization problems. However, a major limitation of QAOA is that it is a “local” algorithm for finite circuit depths, meaning it can only optimize over local properties of the graph. In this paper, we present Phantom-QAOA, a new QAOA ansatz that introduces only one additional parameter to the standard ansatz—regardless of system size—allowing QAOA to “see” more of the graph at a given depth $p$. We achieve this by modifying the target graph to include additional $\alpha$-weighted edges, with $\alpha$ serving as a tunable parameter. This modified graph is then used to construct the phase operator and allows QAOA to explore a wider range of the graph’s features. We derive a general formula for our new ansatz at $p=1$ and analytically show an improvement in the approximation ratio for cycle graphs. We also provide numerical experiments that demonstrate significant improvements in the approximation ratio for the Max-Cut problem over the standard QAOA ansatz for $p=1$ and $p=2$ on random regular graphs up to 16 nodes.
@@ I Introduction
Over the past several years, researchers have studied the potential ability of quantum computers to outperform classical computers on a variety of problems and tasks. One such category of problems are combinatorial optimization problems, where – given a finite discrete set of candidate solutions – the objective is to find an optimal solution with respect to some cost function. Such problems appear in many different domains including power systems [20], logistics [17], VLSI [2], and more. Our work focuses on the well-studied Max-Cut problem, which is to determine an optimal partition of the vertices of the graph into two groups such that the number of edges between the two groups is maximized; however, the main ideas and techniques in our work can also be used for other combinatorial optimization problems.
@@
Since the cost Hamiltonian typically encodes the cost function of the underlying combinatorial optimization problem, it is not clear a priori what alternatives one should consider. In particular, with the exception of an affine shift in the cost function, any other cost Hamiltonian will most likely break QAOA’s connection to Quantum Adiabatic Computing (QAC); this connection ensures that QAOA (with optimal parameters) approaches an optimal solution as the number of circuit layers $p$ tends to infinity. In this work, we consider an alternative cost Hamiltonian that is formed by a modified graph, which is constructed by starting with the original input graph and adding additional $\alpha$-weighted edges, with $\alpha$ being a tunable parameter. If one allows optimization of the $\alpha$-parameter, then this QAOA variant does at least as well as standard QAOA, as the standard QAOA is recovered at $\alpha=0$. We call our new QAOA variant Phantom-QAOA.
@@
Similar to Multi-Angle QAOA—and a recent generalization, eXpressive QAOA [22], which adds an extra Pauli-Y component in the mixer Hamiltonian—our method introduces additional parameters; however, Phantom-QAOA only adds a single parameter, regardless of problem size or number of QAOA layers. On the other hand, Multi-Angle QAOA adds an additional $p(n+m-2)$ parameters where $m$ is the number of terms in the original cost Hamiltonian; the addition of such parameters significantly increases the parameter search space, thus making the cost landscape increasingly difficult and computationally intensive to search over.
In this work, we demonstrate theoretical results for single-layer QAOA with our approach. For any way of adding $\alpha$-weighted edges, we provide a general formula for the expectation of an edge being cut for single-layer QAOA. Using this formula, we demonstrate an example where our approach does strictly better than standard QAOA; more specifically, for even-length cycles, we show that, when placing the additional edges in $G$ to form triangles, our ansatz yields an approximation ratio of $0.7925$ whereas standard QAOA only has an approximation ratio of $0.75$.
@@ II Background
II-A Max-Cut Problem
The problem we focus on in this paper is the Max-Cut problem—a canonical NP-hard combinatorial optimization problem [25]. The problem is defined as follows: given an undirected graph $G=(V,E)$, partition the vertex set $V$ into two disjoint subsets $S$ and $V\setminus S$ in such a way that the number of edges with one endpoint in $S$ and the other in $V\setminus S$ is maximized, i.e.,
$\max_{S\subset V}\;\left|\{(u,v)\in E:u\in S,\;v\in V\setminus S\}\right|.$
The best known classical approach to approximating this optimum is the Goemans-Williamson algorithm [11], which employs semidefinite programming coupled with randomized rounding to achieve a worst-case approximation ratio of 0.878.
An equivalent formulation assigns to each vertex $i\in V$ a binary variable $z_{i}\in\{0,1\}$, so that a candidate solution is represented by the vector $z=(z_{1},z_{2},\dots,z_{n})$. The corresponding cut value is given by
$C(z)=\frac{1}{2}\sum_{(u,v)\in E}\Bigl{(}1-(-1)^{z_{u}+z_{v}}\Bigr{)}.$
Alternatively, we can reformulate the above cost function in terms of computational basis states by using the standard Pauli-$Z$ operator representation (with $Z\ket{0}=\ket{0}$ and $Z\ket{1}=-\ket{1}$). The cost function can then be expressed as
$C=\frac{1}{2}\sum_{(u,v)\in E}\Bigl{(}1-Z_{u}Z_{v}\Bigr{)}.$
II-B Quantum Approximate Optimization Algorithm
QAOA is a hybrid quantum-classical algorithm in which a parameterized quantum circuit generates sample solutions, and a classical optimizer varies the circuit parameters in order to maximize (or minimize) the objective function. The QAOA ansatz is inspired by adiabatic quantum computing [1]. The mixer operator $U(B,\beta_{i})$ and the phase operator $U(C,\gamma_{i})$ are applied in an alternating sequence, and the $\beta_{i}$ and $\gamma_{i}$ are the respective parameters for the $i$th layer. The number of layers, $p$, is called the circuit depth.
@@ II-B Quantum Approximate Optimization Algorithm
The phase operator typically encodes the problem we wish to solve. For the Max-Cut problem, the phase operator is generated by the cost Hamiltonian $C$ that represents the cut value of a given partition of the graph’s vertices. In our formulation, each vertex $i$ in the graph $G=(V,E)$ is associated with a binary variable $z_{i}\in\{0,1\}$ so that a computational basis state $|z_{1}z_{2}\cdots z_{n}\rangle$ corresponds to a particular partition of the vertices—for example, with $z_{i}=0$ indicating that vertex $i$ belongs to one subset and $z_{i}=1$ indicating membership in the other. The phase operator is then given by:
@@
The variational quantum state after $p$ layers is
@@
where |$\ket{s}$ is the initial state. In this work, we follow the standard practice of initializing the state in the most excited state of the mixer Hamiltonian: |$\ket{+}^{\otimes n}$. Note that this initial state is the equal superposition of all $2^{n}$ possible cuts in the graph. Equivalently, measuring |$\ket{s}$ would cause each vertex to independently have equal probability of being on either side of the cut; one can show that this implies that each edge is cut with probability $1/2$ and hence measuring |$\ket{s}$ would yield a cut value of $1/2$ the number of edges.
@@
Maximizing the expectation value of $\langle C\rangle$ with respect to the parameters $\boldsymbol{\gamma}$, $\boldsymbol{\beta}$ is the objective of QAOA. Let $F_{p}$ represent the expectation value of the cost Hamiltonian after $p$ layers. The maximum achievable expectation value, $\text{max}_{\boldsymbol{\gamma},\boldsymbol{\beta}}F_{p}(\gamma,\beta)$, increases with $p$. In the limit as $p\to\infty$, QAOA (with optimal parameters) is guaranteed to achieve an approximation ratio of $1$ [10].
$u$$v$$w$$x$$y$
Figure 1: Visualization of graph $G^{\prime}$, where the black lines are the edges from the original graph $G$ and additional $\alpha$-weighted edges are denoted by red lines. For the edge $(xy)$, $d=2$ and $d^{\prime}=1$ denote two existing and one phantom neighbor of $x$, respectively. Similarly, for vertex $y$, $e=1$ and $e^{\prime}=2$. The triangles formed around this edge include: one with vertex $u$ ($f=1$), one with vertex $w$ ($f^{\prime}=1$), and another with vertex $v$ ($f^{\prime\prime}=1$).
II-C Graph Visibility in QAOA
@@
To demonstrate QAOA’s locality, let’s evaluate a single edge term $C_{uv}=\frac{1}{2}(I-Z_{u}Z_{v})$ in the cost Hamiltonian (recall that $C=\sum_{uv}C_{uv}$). The expected cut value of the edge $(u,v)$ at depth $p=1$ is
@@
However, adjusting the cost Hamiltonian that is used in the phase operatory $U(C,\gamma)$ can lead to changes in how much the QAOA can ”see” for a given $p$. One example of the application of a modified phase operator is shown in [24]. We discuss our new phase operator in detail in the next section.
III Phase Operator with Phantom Edges
Mathematically, we introduce a new phase operator $U(\tilde{C},\gamma)$, where $\tilde{C}$ is the Hamiltonian including additional small-weight ”phantom” edges:
@@ III Phase Operator with Phantom Edges
Here, $C$ is the original Max-Cut cost Hamiltonian, $C^{\prime}=\frac{1}{2}\sum_{(u,v)\in E^{\prime}}\Bigl{(}1-Z_{u}Z_{v}\Bigr{)}$ represents the cut value for the additional edges, and $\alpha$ is the weight of the additional edges. The modified phase operator is then:
@@
III-A Methods for Additional Weighted-Edge Placement
In this paper, we present two methods for constructing the augmented graph $G^{\prime}$, which is used to define the modified cost Hamiltonian $\tilde{C}$ that generates the phase operator $U(\tilde{C},\gamma)$ in our QAOA ansatz. These approaches are referred to as the ”Full” and ”Triangle” methods.
The Full method consists of adding $\alpha$-weighted edges between all pairs of vertices in $G$ that do not already share an edge, forming a fully connected graph $G^{\prime}$. In other words, the set of edges that are added are simply the edges of the complement of $G$. Mathematically, if $E$ is the set of edges in the original graph $G$, the new set of edges in $G^{\prime}$, denoted by $E^{\prime}$, includes all possible edges between vertices:
@@ III-A Methods for Additional Weighted-Edge Placement
The additional edges in $G^{\prime}$ are then weighted by the factor $\alpha$ in the new cost Hamiltonian $C^{\prime}$.
The Triangle method adds $\alpha$-weighted edges only between vertices that are two hops away in $G$, forming triangles with the existing edges. Mathematically, for any vertex pair $(u,v)$ such that the shortest path between $u$ and $v$ in $G$ is exactly two edges, we add an edge $(u,v)$ to $G^{\prime}$:
@@
IV Derivation of Max-Cut Expectation Value for Phantom-QAOA
The formula for the unweighted Max-Cut problem for the original QAOA ansatz at $p=1$ (QAOA1) is contained in the following Theorem [12]:
@@ IV Derivation of Max-Cut Expectation Value for Phantom-QAOA
Consider the QAOA1 state $|\gamma,\beta\rangle$ for Max-Cut on a graph $G$.
  • For each edge $(uv)$,
    $\langle\gamma,\beta|C_{uv}|\gamma,\beta\rangle=\frac{1}{2}+\frac{1}{4}\sin(4\beta)\sin\gamma(\cos^{d}\gamma+\cos^{e}\gamma)$
    $-\frac{1}{4}\sin^{2}(2\beta)\cos^{d+e-2f}\gamma(1-\cos^{f}(2\gamma))=:\langle C_{uv}\rangle(d,e,f)$
    where $d=\deg(u)-1$, $e=\deg(v)-1$, and $f$ is the number of triangles in the graph containing $(uv)$.
  • The overall expectation value is
    $\langle C\rangle=\langle\gamma,\beta|C|\gamma,\beta\rangle=\sum_{(d,e,f)}\langle C_{uv}\rangle(d,e,f)\chi(d,e,f),$
    where $\chi(d,e,f)$ gives the number of edges in $G$ with neighborhood parameters $(d,e,f)$.
In this section, we generalize Theorem 1 for Phantom-QAOA at $p=1$. We acknowledge that the expression itself may not provide immediate insights or be easily interpretable. We provide an example in the following subsection (Section  IV-A) to improve interpretability as well as show the utility of introducing additional edges so that additional triangles are formed with the edges in $G$. For additional clarity on the parameters of the following theorem, we include an example graph with labeled parameters in Fig.  1.
Theorem 2.
For each vertex $u\in V$, define its neighbor sets as $N(u)=\{w\in V:(uw)\in E\}\quad$ and $N^{\prime}(u)=\{w\in V:(uw)\in E^{\prime}\setminus E\}$. Then, for the single-layer Phantom-QAOA algorithm with parameters $\beta$,$\gamma$ and $\alpha$, the expected Max-Cut for each edge $(uv)\in E$ is given by
$\langle C_{uv}\rangle$ $=\frac{1}{2}+\frac{\sin(4\beta)}{4}\sin(\gamma)\bigg{(}(\cos^{d}(\gamma)\cos^{d^{\prime}}(\alpha\gamma)$
$\quad\quad\quad\quad\quad+\cos^{e}(\gamma)\cos^{e^{\prime}}(\alpha\gamma)\bigg{)}$
@@
$\quad\quad\cos^{f^{\prime}}(\gamma(1-\alpha))-\cos^{f^{\prime}}(\gamma(1+\alpha))\bigg{)}$
@@
$\langle\gamma,\beta|C_{uv}|\gamma,\beta\rangle=\frac{1}{2}-\frac{1}{2}\langle s|e^{i\gamma\tilde{C}}e^{i\beta B}Z_{u}Z_{v}e^{-i\beta B}e^{-i\gamma\tilde{C}}|s\rangle.$ (3)
@@
$\langle Z_{u}Z_{v}\rangle$
$=\langle s|e^{i\gamma\tilde{C}}\left(c^{2}Z_{u}Z_{v}+sc(Y_{u}Z_{v}+Z_{u}Y_{v})+s^{2}Y_{u}Y_{v}\right)e^{-i\gamma\tilde{C}}|s\rangle.$ (4)
@@
$=\langle s|(\cos(\gamma)I-i\sin(\gamma)Z_{u}Z_{v})\prod_{i=1}^{d}(\cos(\gamma)I-i\sin(\gamma)Z_{u}Z_{w_{i}})$
×$\quad\times\prod_{i=1}^{d^{\prime}}(\cos(\alpha\gamma)I-i\sin(\alpha\gamma)Z_{u}Z_{w_{i}^{\prime}})|s\rangle,$ (5)
where $d=|N(u)\setminus v|$, $w_{i}$ is a vertex in $N(u)\setminus v$, $d^{\prime}=|N^{\prime}(u)|$, and $w^{\prime}_{i}$ is a vertex in $N^{\prime}(u)$. The only term that contributes to the expectation value in Eq. IV is the one proportional to $Z_{u}Z_{v}\otimes I^{\otimes d}\otimes I^{\otimes d^{\prime}}\cdot Y_{u}Z_{v}=-iX_{u}$. Thus, we have:
@@
Due to symmetry, the expression for the expectation value of the term $Z_{u}Y_{v}$ will be the same as in Eq. IV, except now we use the variables $e=|N(v)\setminus u|$ and $e^{\prime}=|N^{\prime}(v)|$. Thus,
@@
$=\sin\gamma\left(\cos^{d}\gamma\cos^{d^{\prime}}\alpha\gamma+\cos^{e}\gamma\cos^{e^{\prime}}\alpha\gamma\right).$ (7)
For the final term in Eq.  IV, we can rewrite it as
$\langle s|e^{i\gamma\tilde{C}}Y_{u}Y_{v}e^{-i\gamma\tilde{C}}|s\rangle=\langle s|\prod_{i=1}^{d}(\cos(\gamma)I-i\sin(\gamma)Z_{u}Z_{w_{i}})$
@@
×$\quad\times\prod_{j=1}^{e^{\prime}}(\cos(\alpha\gamma)I-i\sin(\alpha\gamma)Z_{v}Z_{w^{\prime}_{j}})Y_{u}Y_{v}|s\rangle.$ (8)
The analysis is similar to the one detailed in [12] for deriving Eq. 5.14, but now we must consider the following three cases for a pair of vertices $(u,v)\in E$:
@@
    Both edges $(u,w_{i})$ and $(v,w_{i})$ are contained in $G$.
@@
    One of the edges, either $(u,w_{i})$ or $(v,w_{i})$, is contained in $G$, while the other edge is contained in $G^{\prime}\setminus G$.
@@
    Both edges $(u,w_{i})$ and $(v,w_{i})$ are contained in $G^{\prime}\setminus G$.
The lowest-order contributions from Case $1$ arise from terms proportional to
$I^{\otimes(d+e-2)}\otimes I^{\otimes(d^{\prime}+e^{\prime})}\otimes(Z_{u}Z_{w})\otimes(Z_{v}Z_{w}),$
with each such term corresponding to one triangle; there are $f$ triangles of this type. Since for any given neighbor $w_{i}$ we have $Z_{u}Z_{w_{i}}\cdot Z_{v}Z_{w_{i}}=I$, any contribution involving an even number of these factors cancels out. Consequently, only terms with an odd number of sine factors (i.e., odd-order terms) contribute. The next nonzero contribution arises at third order, involving three pairs $(Z_{u}Z_{w_{i}},Z_{v}Z_{w_{i}})$; there are $\binom{f}{3}$ such terms, each proportional to $s^{\prime 6}$. This pattern continues for all higher odd orders, and we get
$\binom{f}{1}c^{\prime d+e-2}c^{\prime\prime d+e}s^{\prime 2}+\binom{f}{3}c^{\prime d+e-6}c^{\prime\prime d+e}s^{\prime 6}+$
$\binom{f}{5}c^{\prime d+e-10}c^{\prime\prime d+e}s^{\prime 10}+\ldots$
Similarly, in Case 2, the first-order contributions are proportional to
$I^{\otimes(d+e-1)}\otimes I^{\otimes(d^{\prime}+e^{\prime}-1)}\otimes(Z_{u}Z_{w})\otimes(Z_{v}Z_{w^{\prime}}),$
corresponding to $f^{\prime}$ triangles. In Case 3, the first-order contributions are proportional to
$I^{\otimes(d+e)}\otimes I^{\otimes(d^{\prime}+e^{\prime}-2)}\otimes(Z_{u}Z_{w^{\prime}})\otimes(Z_{v}Z_{w^{\prime}}),$
with $f^{\prime\prime}$ triangles. As in Case 1, only odd-order terms contribute, and the higher-order contributions in both cases follow the same combinatorial pattern. We thus obtain the following:
@@
Applying the binomial theorem and simplifying via trigonometric identities leads to the following expression:
@@
$=c^{d+e-2f}c^{\prime\prime d^{\prime}+e^{\prime}}(1-\cos^{f}(2\gamma))$
$+c^{d+e-f^{\prime}}c^{\prime\prime d^{\prime}+e^{\prime}-f^{\prime}}\frac{1}{2}\left(\cos^{f^{\prime}}(\gamma(1-\alpha))-\cos^{f^{\prime}}(\gamma(1+\alpha)\right)$
$+c^{d+e}c^{\prime\prime d^{\prime}+e^{\prime}-2f^{\prime\prime}}(1-\cos^{f^{\prime\prime}}(2\alpha\gamma)).$ (9)
Substituting Eq.  IV and Eq.  IV into Eq.  IV we obtain the final expression for $\langle C_{uv}\rangle$ in terms of $\gamma,\beta$ and $\alpha$.